OpenStudy (anonymous):

a2. A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 1. Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. 2. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). 3. Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

OpenStudy (bibby):

Try multipling the first equation by 2. Then add it to the second equation. Solve (substitution or elimination or whatever other methods you can think of)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how would i add them together.

OpenStudy (bibby):

When adding equations you add like terms. For example 2x+3y + 4x + 8y = 6x + 11y

OpenStudy (anonymous):

ok so in this case its, since its times 2 , 4x + 14y = 2?

OpenStudy (bibby):

That's right. Then we add it to -3x – 4y = 5

OpenStudy (anonymous):

then it would be 1x + 18y = 7?

OpenStudy (bibby):

14y+(-4y)= 10y

OpenStudy (bibby):

by that I mean 14y + (-4y) = 14y-4y = 10y

OpenStudy (anonymous):

ok thanks alot it really helped!!

OpenStudy (bibby):

 but we aren't done